OpenStudy (anonymous):

what is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y =-1

OpenStudy (anonymous):

-7 7 9 -9

OpenStudy (anonymous):

the options

OpenStudy (jack1):

simultaneous equations: http://www.bbc.co.uk/schools/gcsebitesize/maths/algebra/simultaneoushirev1.shtml

OpenStudy (jack1):

do you understand how they work?

OpenStudy (anonymous):

@kelliegirl33 thats not in the options so idk

OpenStudy (anonymous):

I messed up...let me do it again

OpenStudy (anonymous):

5x + 4y = 1 -->(4)5x + 4y = 1 4x + 3y = - 1-->(-5)4x + 3y = - 1 ----------------- 20x + 16y = 4 (result of multiplying by 4) -20x - 15y = 5 (result of multiplying by -5) -----------------add 0 + y = 9 y = 9

OpenStudy (anonymous):

thank you can you help me with another?

OpenStudy (anonymous):

possible...let me see what you got

OpenStudy (anonymous):

auditions for a play in New York were held with 120 men and women trying out for the lead singer position. if twice the number of men less three times the number of women id equal to 20, how many women auditioned?

OpenStudy (anonymous):

m + w = 120 --> m = 120 - w 2m - 3w = 20 now sub 120 - w in for m 2(120 - w) - 3w = 20 240 - 2w - 3w = 20 240 - 5w = 20 240 - 20 = 5w 220 = 5w 220/5 = w 44 = w 44 women auditioned You might need to get a second opinion because I am not sure if I got the second equation right.

OpenStudy (anonymous):

@UnkleRhaukus ....is the second equation correct ?

OpenStudy (anonymous):

The wording for the second equation is confusing me

OpenStudy (anonymous):

@skullpatrol ...is the second equation correct ?

OpenStudy (anonymous):

maybe somebody will come and check this

OpenStudy (anonymous):

thats there as on option so i think its right

OpenStudy (anonymous):

I wasn't sure if the second equation was 2m - 3w = 20 or 3w - 2m = 20 ?

OpenStudy (anonymous):

@terenzreignz ....help please

terenzreignz (terenzreignz):

you can "cheat"... or you can solve this normally, as if there were no choices... what's your preferred method?

OpenStudy (unklerhaukus):

think you have interpreted the grama as i would have @kelliegirl33

OpenStudy (anonymous):

If I use the other equation....the 3w - 2m = 20....the answer does not come out even, so I think my answer is correct